# DDOS: The Drone Depth and Obstacle Segmentation Dataset

### v1

@@ Abstract
Accurate depth and semantic segmentation are crucial for various computer vision tasks. However, the scarcity of annotated real-world aerial datasets poses a significant challenge for training and evaluating robust models. Additionally, the detection and segmentation of thin objects, such as wires, cables, and fences, present a critical concern for ensuring the safe operation of drones. To address these limitations, we present a novel synthetic dataset specifically designed for depth and semantic segmentation tasks in aerial views. Leveraging photo-realistic rendering techniques, our dataset provides a valuable resource for training models using a synthetic-supervision training scheme while introducing new drone-specific metrics for depth accuracy.
@@ 1 Introduction
The safe operation of drones typically requires the accurate detection and segmentation of thin objects, such as wires, cables, and fences, which can pose as dangerous obstacles. Detecting and segmenting these objects with precision is crucial for enabling efficient drone navigation and avoiding potential hazards. However, due to the fine spatial characteristics and limited representation of thin objects in existing datasets, this domain remains relatively understudied in aerial analysis [11].
Existing datasets for depth estimation and semantic segmentation in aerial imagery primarily focus on common objects like buildings, vehicles, and trees [4, 15]. While these datasets have facilitated significant progress in aerial analysis, they do not adequately address the unique challenges posed by thin objects. Thin objects often exhibit intricate and elongated structures, making them difficult to detect and accurately segment from aerial views. Moreover, the limited occurrence of thin objects in existing datasets hampers the development and evaluation of robust algorithms.
To bridge this gap, we propose the Drone Depth and Obstacle Segmentation (DDOS) dataset, a synthetic dataset specifically designed to address the scarcity of labeled real-world data and emphasize the detection and segmentation of thin objects from aerial perspectives. DDOS leverages advanced techniques in computer graphics and rendering to generate synthetic aerial images that closely resemble real-world scenes. By incorporating photo-realistic qualities, DDOS enables the exploration of thin object detection and segmentation algorithms in a controlled environment.
The synthetic nature of DDOS offers several advantages over traditional real-world datasets. First, it provides a substantially larger and diverse set of labeled examples of thin objects, which are typically scarce in real-world scenarios. This abundance of labeled data enables more effective training and evaluation of thin object detection and segmentation algorithms. Second, by generating synthetic images, we have access to pixel-perfect ground truth annotations, depth, optical flow and surface normals, which allows for a cleaner training signal.
USF [2] | NE-VBWD [13] | Mid-Air [4] | TartanAir [15] | DDOS (ours) | |
Data type | Real | Real | Synthetic | Synthetic | Synthetic |
Flight Trajectories | |||||
Frames | † | ||||
Labeled frames | 91 | † | |||
Resolution | |||||
Frame rate | - | ||||
Environment | Town | Town/Nature | Nature | Various | Town/Nature |
Camera motion | Handheld | Helicopter | Drone | Random | Drone |
Camera pose | No | No | Yes | Yes | Yes |
Optical flow | No | No | No | Yes | Yes |
Depth map | No | Sparse | Dense | Dense | Dense |
Labeled segmentation | Wires only | Wires only | Yes | No* | Yes |
Thin objects | Yes | Yes | No | No* | Yes |
Mesh objects | No | No | No | No* | Yes |
To facilitate a comprehensive understanding of the dataset, we conduct a detailed statistical analysis of DDOS, focusing on various key attributes such as class density, flight characteristics and spatial distribution. Additionally, we compare our dataset with existing datasets for depth estimation and semantic segmentation in aerial imagery. This comparative analysis aims to underscore the distinctive contributions of DDOS in addressing the challenges posed by thin objects, providing valuable insights into its potential applications and advantages over current datasets in the setting of aerial analysis.
In addition, we introduce new drone-specific metrics tailored specifically to assess depth accuracy in the context of thin object detection and segmentation. These metrics aim to provide a more nuanced and relevant evaluation framework for algorithms operating in drone applications. By accounting for the distinctive challenges posed by thin objects, these metrics offer a more accurate reflection of performance in real-world scenarios, thereby contributing to the advancement of drone technology and safety.
@@
Finally, we present baseline results obtained by applying state-of-the-art algorithms to the DDOS dataset, offering a benchmark for future research in thin object detection and segmentation. The discussion of these baseline results delves into the strengths and limitations of existing approaches, highlighting opportunities for refinement and innovation.
@@
DDOS Dataset: We present the Drone Depth and Obstacle Segmentation (DDOS) dataset, specifically designed to tackle the intricate task of detecting and segmenting thin objects.
Statistical Analysis and Dataset Comparison: A comprehensive statistical analysis of DDOS is conducted, shedding light on various key attributes, including object density, spatial distribution, and structural complexity. This analysis is complemented by a meticulous comparison of DDOS against relevant datasets in the field.
@@
Baseline Results and Discussion: Baseline results, achieved through state-of-the-art algorithms applied to the DDOS dataset, are presented. These results serve as a benchmark for future research in thin object detection and segmentation. We engage in a thorough discussion, exploring in-depth insights into the strengths and limitations of the proposed metrics and shedding light on potential avenues for further research and improvement.
@@ 2 Related Work
The availability of comprehensive and high-quality datasets for thin object detection is limited.
@@ 2.1 Driving datasets
Prominent datasets such as KITTI [6, 10] and Cityscapes [3], while widely utilized in computer vision, are not optimally suited for the specific challenges posed by drone applications.
KITTI, a benchmark dataset for autonomous driving, offers rich annotations for object detection, tracking, and scene understanding. However, its primary focus on road scenes and vehicular scenarios does not align with the nuances of drone navigation. The dataset lacks diversity in the types of thin objects encountered by drones, particularly in scenarios involving cables, wires, and similar structures commonly found in aerial environments. Additionally, KITTI’s distance range and object scale may not fully encompass the variability inherent in drone perspectives, limiting its applicability for developing robust thin object detection models for drone operations.
Similarly, Cityscapes, designed for urban scene understanding, provides detailed semantic segmentation annotations but falls short in addressing the specific challenges faced by drones. In particular, thin objects like wires and cables are not adequately represented. Consequently, models trained on KITTI or Cityscapes may exhibit limitations when applied to drone scenarios, where the detection of slender structures is crucial for safe navigation.
@@ 2.2 Wire detection datasets
Some datasets focus solely on wire segmentation, such as USF [2] and NE-VBWD [13]. The USF dataset, a widely used resource, contains just under images of relatively low quality, with only around images featuring wires accompanied by labeled annotations. Notably, the dataset lacks depth information, and the provided ground truth wire annotations are not pixel-accurate. Additionally, wire labels are often either missing or incomplete, as the dataset defines all wires as straight lines, neglecting their potential curvature. This limitation is significant given that wires in real-world scenarios can exhibit substantial curves. Addressing these issues is crucial for advancing the accuracy and applicability of thin object detection models.
NE-VBWD [13] represents a more recent dataset, albeit small, offering both pixel-wise annotations and information regarding the distance to the wires. Notably, this dataset places emphasis on long-range wire detection, making it better suited for manned aircraft rather than drones. Specifically, the majority of wires in the NE-VBWD dataset are situated at a distance of approximately , with very few within . This skew towards longer distances is attributed to the dataset’s acquisition via a helicopter, which, due to safety considerations, could not closely approach the wires. Consequently, NE-VBWD proves less useful for drone applications, where the primary concern lies in the detection of wires within the range of –, while those beyond this range are of comparatively lesser importance.
RGB Image | Depth Map | Segmentation | Optical Flow | Surface Normals |
@@ 2.3 Drone datasets
Both Mid-Air [4] and TartanAir [15] are synthetic datasets generated using Airsim [12]. Mid-Air offers high-resolution images at a high frame rate, with a focus on nature scenes and limited inclusion of man-made objects. In contrast, TartanAir provides numerous trajectories spanning various environments such as towns, cities, nature, indoor settings, and even underwater. However, TartanAir’s segmentation is hindered by assigning objects to random classes, and variations of the same object are placed into different classes. This unhelpful segmentation diminishes the utility of TartanAir for a comprehensive training platform.
Our dataset represents a significant advancement, providing high-resolution images that encompass a rich array of challenging semantic segmentation classes. This diversity spans both intricate man-made environments and the nuanced complexity of natural scenes. By incorporating a broad spectrum of scenarios, our dataset offers a more comprehensive and realistic training platform for algorithms to navigate and interpret varied environments.
To underscore the distinctiveness of our dataset, we present a detailed comparison with other relevant datasets in Table 1. This comparative analysis aims to highlight the unique attributes and advantages of our dataset over existing alternatives, emphasizing its suitability for addressing the intricate task of thin object detection and segmentation in a wide range of real-world scenarios.
@@ 3 Dataset Features
We introduce the Drone Depth and Obstacle Segmentation (DDOS) dataset, a synthetic benchmark explicitly designed to tackle the scarcity of annotated real-world data and cater to the challenges of obstacle segmentation and depth estimation from drone perspectives. Comprising unique and diverse drone flights conducted in simulated environments, the DDOS dataset offers a photo-realistic representation of real-world scenarios encountered by drones. This dataset serves as a crucial resource for advancing the field, addressing the limitations associated with the lack of annotated real-world data.
Focusing specifically on thin objects that present significant safety concerns for drone flights, such as cables and wires, DDOS emerges as a valuable resource for the development of algorithms capable of precise segmentation and depth estimation in intricate environments. Each flight in the dataset captures high-resolution images utilizing a monocular camera affixed to the front of the drone. Accompanying these images are corresponding depth maps, pixel-level object segmentation masks, optical flow information, and surface-normals. All image modalities maintain a resolution of , and the depth maps cover a range from to .
Carefully curated, DDOS is partitioned into training, validation, and testing subsets, encompassing 300, 20, and 20 flights, respectively. Figure 1 provides visual examples, illustrating the diverse classes represented in the dataset, with additional examples shown in Appendix B. This meticulous curation and emphasis on thin objects contribute to the dataset’s efficacy in addressing the intricacies of obstacle segmentation and depth estimation relevant to drone navigation and safety.
DDOS features pixel-wise object segmentation masks with ten distinct classes, allowing for detailed analysis of diverse obstacles and environmental elements. These classes, namely ultra thin, thin, small mesh, large mesh, trees, buildings, vehicles, animals, other, and background, provide a comprehensive breakdown. For instance, the ultra thin class covers objects like wires and cables, while the thin class encompasses streetlights and poles. The small mesh class includes objects like fences and nets, and the large mesh class involves structures similar to pylons and radio masts. This comprehensive categorization within DDOS serves as a robust resource, facilitating the development and assessment of advanced algorithms specifically tailored to tackle the complex challenges of obstacle segmentation and depth estimation from the unique vantage point of drones.
The synthetic nature of the DDOS dataset overcomes the limitations of scarce annotated real-world data, allowing researchers to systematically explore and compare different approaches in a controlled environment. By leveraging this dataset, researchers can advance the field of drone vision, developing innovative solutions to address the challenges of obstacle segmentation and depth estimation. Ultimately, DDOS aims to enhance the safety and efficiency of drone operations by facilitating the development of robust algorithms capable of accurately detecting and navigating around thin objects, such as cables and wires, which are critical for ensuring the safety of drone flights.
@@ 4 Data Generation
The images in DDOS are generated using AirSim[12], a drone simulator. DDOS is composed of two environments that mimic real-world scenarios. The first environment resembles a small suburban town, featuring dense trees and numerous power lines, replicating the challenges faced during drone flights in residential areas. The second environment represents a park setting, incorporating elements such as a football field with floodlights, a beach volleyball court, dense trees as well as office buildings. These environments collectively offer diverse obstacles and structures, allowing researchers to develop and evaluate algorithms capable of addressing the complexities associated with different real-world environments. By encompassing characteristics like dense tree coverage, power lines and varying weather conditions, such as snow cover, the dataset provides a comprehensive platform for advancing obstacle segmentation and depth estimation algorithms for safe and effective drone flights.
@@
In order to promote relatively safe flight paths, we developed a dynamic obstacle detection algorithm to modify intermediate targets in response to potential collision risks. This algorithm utilizes the most recent ground truth depth map obtained during the recorded flight observations. By empirically determining a threshold, objects that are deemed too close trigger updates to the intermediate targets. The updated targets are strategically adjusted based on the detected obstacle’s location, causing the drone to navigate away from the identified collision risk. This obstacle avoidance approach is not flawless, especially when dealing with thin objects, occasional collisions resulting in crashes still occur. In such cases, the observations associated with the crash event are discarded, and the flight process is restarted to ensure data integrity. It is important to note, the collision avoidance mechanism is purposefully designed to be lax, as near misses and even minor crashes can offer valuable data points for training purposes.
@@
To uphold the overall integrity of the dataset and exclude instances of undesired behavior, additional validation criteria are applied after flight generation. These criteria serve to filter out scenarios where the drone becomes stuck or encounters unusual situations, such as becoming entangled in trees. By incorporating these post-flight validation steps, the dataset ensures that the collected observations reflect reliable and meaningful flight behaviors, enabling robust algorithm training and evaluation.
@@
Incorporating diverse weather conditions, such as snow cover and wet roads, is crucial for enhancing the robustness of computer vision models, particularly in the context of autonomous systems and transportation. Adverse weather introduces challenges like reduced visibility and altered surface characteristics, which significantly impact the performance of vision-based algorithms. Snow cover, for instance, introduces high reflectivity and obscures road markings, demanding algorithms capable of adapting to these dynamic conditions. Similarly, wet roads can lead to glare and reflections, posing challenges for accurate object detection and scene understanding. By including such weather scenarios in datasets, researchers can develop and evaluate models that demonstrate resilience in real-world, dynamic environments, ultimately advancing the reliability of computer vision systems across various weather conditions.

### v2

@@ Abstract
The advancement of autonomous drones, essential for sectors such as remote sensing and emergency services, is hindered by the absence of training datasets that fully capture the environmental challenges present in real-world scenarios, particularly operations in non-optimal weather conditions and the detection of thin structures like wires. We present the Drone Depth and Obstacle Segmentation (DDOS) dataset to fill this critical gap with a collection of synthetic aerial images, created to provide comprehensive training samples for semantic segmentation and depth estimation. Specifically designed to enhance the identification of thin structures, DDOS allows drones to navigate a wide range of weather conditions, significantly elevating drone training and operational safety. Additionally, this work introduces innovative drone-specific metrics aimed at refining the evaluation of algorithms in depth estimation, with a focus on thin structure detection. These contributions not only pave the way for substantial improvements in autonomous drone technology but also set a new benchmark for future research, opening avenues for further advancements in drone navigation and safety.
@@ 1 Introduction
Fully autonomous drones are poised to revolutionize a multitude of sectors, including remote sensing [35, 16, 26, 17, 3, 31], package delivery [4, 13], emergency services, and disaster response [10, 2, 11, 28, 9, 29]. While manually controlled drones have been effectively employed in specific sectors, the advent of fully autonomous drones is poised to unlock an array of novel applications, enhancing efficiency and expanding capabilities. However, realizing this potential is contingent upon the ability of drones to navigate safely and autonomously, which in turn requires a precise understanding of their environment. Current datasets for training drone navigation systems are inadequate, particularly in representing challenging scenarios such as the detection of thin structures like wires and cables, and operation under diverse weather conditions [25]. This deficiency highlights the need for a dataset that provides a comprehensive representation of the environment, enabling accurate semantic segmentation and depth estimation across a wide range of objects and conditions.
To address this gap, we introduce the Drone Depth and Obstacle Segmentation (DDOS) dataset, a novel resource designed to significantly enhance the training of autonomous drones. DDOS stands out for its dual emphasis on depth and semantic segmentation annotations, with a particular focus on the precise identification of thin structures (a critical but often overlooked aspect in existing datasets). By incorporating advanced computer graphics and rendering techniques, DDOS generates synthetic aerial images that mirror the complexity of real-world environments, encompassing a variety of settings and weather conditions ranging from clear skies to adverse weather scenarios such as rain, fog, and snowstorms.
USF | NE-VBWD | TTPLA | PIM | UAVid | AeroScapes | Ruralscapes | Mid-Air | TartanAir | SynthWires | SynDrone | DDOS | |
[7] | [33] | [1] | [36] | [21] | [27] | [23] | [12] | [38] | [22] | [30] | (ours) | |
Data type | Real | Real | Real | Real | Real | Real | Real | Synthetic | Synthetic | Synthetic | Synthetic | Synthetic |
Flight Trajectories | NA | |||||||||||
Frames | † | |||||||||||
Labeled frames | 91 | * | † | |||||||||
Resolution | ||||||||||||
Frame rate | - | - | - | - | ||||||||
Environment | Town | Town/Nature | Pylons | Pylons | Town/Nature | Various | Town/Nature | Nature | Various | Various | Town | Town/Nature |
Camera motion | Handheld | Helicopter | Drone | Drone | Drone | Drone | Drone | Drone | Random | Drone | Drone | Drone |
Altitude | - | - | - | - | - | - | ||||||
Weather variations | No | No | No | No | No | No | No | Yes | No | No | No | Yes |
Camera pose | No | No | No | No | No | No | No | Yes | Yes | No | Yes | Yes |
Optical flow | No | No | No | No | No | No | No | No | Yes | No | No | Yes |
Depth map | No | Sparse | No | No | No | No | No | Yes | Yes | No | Yes | Yes |
Segmentation | Wires only | Wires only | Yes | No | Yes | Yes | Yes | Yes | No‡ | Wires only | Yes | Yes |
Thin structures | Yes | Yes | Yes | Patches | No | Yes | No | No | No‡ | Yes | No | Yes |
Mesh structures | No | No | Rough | Patches | No | Large only | No | No | No‡ | No | No | Yes |
Our objectives with the DDOS dataset are twofold: firstly, to provide a richly annotated resource that reflects the diversity of scenarios encountered by drones, with a particular focus on thin structures and adverse weather conditions. Secondly, to enable the development and evaluation of algorithms that significantly improve the safety, reliability, and operational efficiency of autonomous drones. By achieving these objectives, we aim to bridge the gap in existing datasets and facilitate the advancement of drone technology to meet the demands of real-world applications.
We present a thorough analysis of DDOS which explores key characteristics including class density, flight dynamics, and spatial distribution, providing a granular understanding of its composition and capabilities. Through comparative analysis with existing datasets, we highlight DDOS’s contributions such as incorporating numerous thin and ultra-thin structures with accurate depth and segmentation labels, as well as diverse weather conditions. Furthermore, we propose new drone-specific metrics designed to accurately evaluate class-specific depth estimation performance. These metrics are tailored to reflect the operational realities of drone applications, offering a refined lens through which to assess algorithmic performance and contributing to the broader goal of advancing drone technology and safety.
@@
Finally, we present baseline results obtained by applying state-of-the-art algorithms to the DDOS dataset, establishing a benchmark for future research in thin object detection. We examine the strengths and limitations of current methodologies, particularly highlighting their notable failure to accurately predict the depth of thin structures. This analysis emphasizes significant opportunities for refinement and innovation within this domain.
@@
DDOS Dataset: We present the Drone Depth and Obstacle Segmentation (DDOS) dataset, a comprehensive resource developed to significantly improve the training of autonomous drones through extensive depth and semantic segmentation annotations, with a special focus on accurately identifying thin structures.
Statistical Analysis and Dataset Comparison: We provide a thorough examination of the DDOS dataset, highlighting its unique attributes such as class distributions, spatial distribution, and flight dynamics. Our analysis is enriched by a detailed comparative study, positioning DDOS in the broader context of existing datasets and underscoring its distinctive value in addressing specific challenges in drone navigation.
@@
Baseline Results and Discussion: We present baseline results from applying state-of-the-art algorithms to the DDOS dataset, establishing benchmarks for thin object detection research. Our discussion identifies a critical shortfall in existing depth estimation methods, emphasizing the need for future advancements.
@@ 2 Related Work
The scarcity of high-quality drone datasets hampers autonomous drone training. This section reviews relevant datasets, evaluating their strengths and weaknesses in regards to training autonomous drones. These evaluations are summarized in Table 1.
@@ 2.1 Driving datasets
The KITTI [15, 24], Cityscapes [8], nuScenes [6], and Waymo [34] datasets, essential in computer vision for autonomous driving, fall short in addressing drone-specific requirements. KITTI’s concentration on road scenes lacks the aerial views and diverse thin structures crucial for drone navigation. Similarly, Cityscapes, nuScenes, and Waymo fail to capture the unique aerial perspectives and the slender objects like wires and cables vital for drone safety. The absence of these aerial viewpoints and the limited representation of thin structures mean that models trained on these datasets are not fully equipped to meet the challenges of drone-based navigation.
@@ 2.2 Wire detection datasets
Several datasets have been specifically designed to tackle the challenge of wire detection, given its critical importance for ensuring the safety of low-flying drones.
The USF dataset [7] and NE-VBWD [33] are pivotal resources dedicated to wire detection, offering a unique perspective on the challenges of identifying thin structures in aerial imagery. The USF dataset, while extensive, is limited by its image quality and the accuracy of its wire annotations, which are not pixel-accurate and often overlook the real-world curvature of wires, instead defining them as straight lines. This simplification fails to capture the complexity of wire shapes in various environments, undermining the dataset’s utility for training models to detect thin structures accurately. NE-VBWD, although a more recent addition, offers pixel-wise annotations and distance information, focusing on long-range wire detection. However, its suitability for drone applications is limited due to its emphasis on wires located at distances more relevant to manned aircraft, thus diminishing its relevance for low-altitude drone operations where proximity to wires is a critical safety concern.
TTPLA [1] and PIM [36] also contribute to the field by focusing on transmission towers and power lines, with TTPLA utilizing drone imagery but lacking depth information, and PIM providing small image patches for wire detection without offering semantic segmentation. These datasets, while enriching the domain with specific insights into wire and tower detection, similarly fall short in addressing the broad needs of autonomous drone navigation, such as a diverse range of thin structures, depth mapping, and environmental conditions beyond the mere presence of wires.
@@ 2.3 Drone datasets
UAVid [21], AeroScapes [27], and Ruralscapes [23] serve as general drone datasets. They provide a broader view of urban and rural landscapes from a drone’s perspective, including various object classes for semantic segmentation. Despite their wider scope, these datasets still lack sufficient emphasis on thin structures, such as wires, which are crucial for the safe navigation of drones in complex environments.
SynthWires [22] utilizes a different approach by overlaying synthetic wires over real-world images from drones. This method enhances the variety of wire scenarios available for training, although the absence of depth information limits the dataset’s applicability for comprehensive 3D navigation and obstacle avoidance training.
In enhancing the dataset landscape for drone navigation research, Mid-Air [12], TartanAir [38], and SynDrone [30] represent significant contributions as synthetic datasets offering voluminous labeled training samples. These datasets play a pivotal role in simulating a diverse array of flight dynamics and environmental conditions, providing essential assets such as precise depth maps and camera poses critical for the advancement of sophisticated drone navigation algorithms. Despite their value, these datasets exhibit certain limitations that restrict their comprehensive utility in fully leveraging the potential of synthetic data generation.
One notable shortfall is their failure to encapsulate a complete spectrum of flight scenarios, particularly those involving close encounters, aggressive maneuvering, and very low-altitude flying. Such scenarios, while perilous to execute in real-world settings, are quintessential for preparing drones to navigate through complex, unpredictable environments. Synthetic datasets, with their capacity for controlled simulation, are uniquely positioned to safely incorporate these high-risk flight patterns, thereby enriching the training regime without endangering equipment or safety.
Moreover, while synthetic datasets offer the advantage of generating pixel-perfect segmentation and precise depth measurements, especially for thin structures – attributes unattainable with conventional data collection methods – they fall short in representing thin structures like wires, cables, and fences. These elements are critical for ensuring the navigational reliability of drones in densely populated or structurally complex areas. The absence of such objects in the datasets underscores a missed opportunity to leverage some of the benefits of synthetic data generation.
Our proposed dataset, DDOS, is designed to surpass the limitations of existing datasets in wire detection and drone navigation. It provides detailed representations of thin structures and a wide array of other entities, incorporating weather variability and extensive drone motion. Its synthetic foundation enables simulations of close encounters with objects, typically unsafe in reality, enhancing the dataset’s utility and realism for drone training.
Image Depth Segmentation Flow Surface normals
@@ 3 Dataset Features
We introduce the DDOS dataset, specifically designed for the training of autonomous drones, utilizing synthetic data generation to compile unique drone flights. This dataset is characterized by its comprehensive coverage of various weather conditions, from clear skies to snowstorms, and includes high-risk scenarios such as close encounters and minor collisions. These scenarios, crucial for drone training, are typically too hazardous to replicate in real-world settings. The dataset is notable for its provision of pixel-level precision in semantic segmentation and depth information, particularly for challenging objects such as wires, cables, and fences, thus offering a photo-realistic simulation of environments drones are likely to encounter.
Each flight within the DDOS dataset consists of frames, culminating in a total of frames across the dataset. This substantial volume of data supports detailed analysis and algorithm training. The dataset emphasizes thin structures, which present significant navigational challenges, thereby serving as a critical resource for the development of algorithms that require precise segmentation and depth estimation capabilities in complex aerial scenarios. Accompanying the high-resolution images captured by a monocular front facing camera are depth maps, semantic segmentation masks, optical flow data, and surface normals. These components are provided at a resolution of pixels, with depth maps covering a range from . Additionally, the dataset incorporates exact drone pose, velocity, and acceleration data for each frame.
The DDOS dataset is systematically divided into training, validation, and testing subsets, consisting of , , and flights, respectively. It features pixel-wise segmentation masks for ten distinct classes, enabling in-depth analysis of various obstacles and environmental elements. Figure 1 displays select examples from the dataset, demonstrating the diversity of classes represented. More examples are available in Appendix B. The methodological approach to dataset generation and the classification scheme are further elaborated in Section 4, providing insight into the dataset’s design choices and structure.
@@ 4 Data Generation
DDOS is generated using AirSim [32], an open-source drone simulator. DDOS is composed of two environments that mimic real-world scenarios. The first environment resembles a small suburban town, featuring dense trees and numerous power lines, replicating the challenges faced during drone flights in residential areas. The second environment represents a park setting, incorporating elements such as a football field with floodlights, a beach volleyball court, dense trees as well as office buildings. These environments collectively offer diverse obstacles and structures, allowing researchers to develop and evaluate algorithms capable of addressing the complexities associated with different real-world environments. By encompassing characteristics like dense tree coverage, power lines, and varying weather conditions, the dataset provides a comprehensive platform for advancing obstacle segmentation and depth estimation algorithms for safe and effective drone flights.
@@
In order to promote relatively safe flight paths, we developed a dynamic obstacle detection algorithm to modify intermediate targets in response to potential collision risks. This algorithm utilizes the most recent ground truth depth map obtained during the recorded flight observations. By empirically determining a threshold, objects that are deemed too close trigger updates to the intermediate targets. The updated targets are strategically adjusted based on the detected obstacle’s location, causing the drone to navigate away from the identified collision risk. This obstacle avoidance approach is not flawless, especially when dealing with thin structures, occasional collisions resulting in crashes still occur. In such cases, the observations associated with the crash event are discarded and the flight process is restarted to ensure data integrity. It is important to note, the collision avoidance mechanism is purposefully designed to be lax, as near misses and even minor crashes can offer valuable data points for training purposes.
@@
To uphold the overall integrity of the dataset and exclude instances of undesired behavior, additional validation criteria are applied after flight generation. These criteria serve to filter out scenarios where the drone becomes stuck or encounters unusual situations, such as becoming entangled in trees. By incorporating these post-flight validation steps, the dataset ensures that the collected observations reflect reliable and meaningful flight behaviors, enabling robust algorithm training, and evaluation.
@@
DDOS encompasses diverse environmental and weather conditions, including sunny, dusk, and brightly lit night scenes, along with rain, fog, snow, and changes due to wet surfaces and snow cover. These conditions challenge vision-based algorithms with reduced visibility and altered surface characteristics, such as increased reflectivity from snow and glare from wet roads, complicating object detection and scene analysis. Including these varied scenarios is essential for developing models that adapt and perform consistently in all real-world settings.
Classes
Objects are systematically classified based on their significance for drone navigation. Ultra Thin encompasses wires and cables; Thin Structures includes poles and signs; Small Mesh pertains to fences and nets; and Large Mesh covers objects such as transmission towers that permit drone passage. Additionally, Trees, Buildings, Vehicles, and Animals are categorized based on straightforward characteristics. The Other class encompasses diverse objects like bus stops, post boxes, chairs, and tables. Background refers to elements such as the ground and sky, providing context within the scene.